OpenStudy (epicteatime):

ompare and contrast the two quadratic equations below. In order to receive full credit, use complete sentences to describe the following: The direction each parabola opens The vertex of each parabola y = x2 - 2x y = -2x2 + 4x - 3

OpenStudy (epicteatime):

@ParthKohli

OpenStudy (epicteatime):

@thomaster

OpenStudy (epicteatime):

YES HELP ME!!!

OpenStudy (epicteatime):

...

OpenStudy (anonymous):

You can tell the direction by looking at a. The equation for this is: Ax + By = C It has been reduced to: By = Ax - c

OpenStudy (anonymous):

If A is negative, the parabola looks like a frown, or opens downwards.

OpenStudy (epicteatime):

no i keep getting (1,-1) for my vortex!!

OpenStudy (epicteatime):

i know how to do it :3

OpenStudy (anonymous):

Let me check. :)

OpenStudy (epicteatime):

k

OpenStudy (epicteatime):

cause its stupid, both of em is (1,-1) for me ;P

OpenStudy (anonymous):

One second..

OpenStudy (epicteatime):

k

OpenStudy (anonymous):

f(x) = x^2 - 2x f(1) = 1^2 - 2(1) f(1) = 1 - 2 f(1) = -1 So I know that that's an ordered pair (maybe not the vertex) f(1) = -2x^2 + 4x - 3 f(1) = -2(1)^2 + 4(1) - 3 f(1) = -2 + 4 - 3 f(1) = -1 So I know that's an ordered pair.

OpenStudy (epicteatime):

sry brb

OpenStudy (anonymous):

I'm gonnna use the quadratic formula for this.

Parth (parthkohli):

1. \(ax^2 + bx + c\) is the general equation of a parabola. If \(a > 0\), then the parabola is an upwards U. If \(a < 0\), then the parabola is a downwards U, which kinda looks like \(\cap\). 2. The vertex of a parabola is given by \(\left(-\dfrac{b}{2a}, f\left(- \dfrac{b}{2a}\right)\right)\). Do you know why that is?

OpenStudy (epicteatime):

kk

OpenStudy (epicteatime):

back

OpenStudy (anonymous):

y = x^2 - 2x \[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac }}{ 2a }\]b = -2 a = 1 c = 0 b = -2\[x = \frac{ 2 \pm \sqrt{-2 ^{2} - 4(1)(0)} }{ 2(1) }\] \[x = \frac{ 2 \pm \sqrt{4 } }{ 2 }\] \[x = \frac{ 2 \pm 2 }{ 2 }\] \[x = \frac{ 2 }{ 2 } \pm \frac{ 2 }{ 2}\] \[x = 1 \pm 1\] \[x = 2, x = 0\] Our x vertex will be between these two points. (1, ?)

OpenStudy (epicteatime):

i got -1

OpenStudy (epicteatime):

TOLD YOU THIS

OpenStudy (anonymous):

You got -1 for y, now we solve for y.

OpenStudy (anonymous):

Then I can see what the vertex is.

OpenStudy (epicteatime):

ITS NOT THAT I DONT KNOW HOW TO DO THIS! ITS THAT I KEEP GETTING 1 AND -1!!!

OpenStudy (anonymous):

I understand that, but what makes you think you're wrong?

OpenStudy (epicteatime):

why! why i keep getting -1!!

OpenStudy (anonymous):

Because it could be the correct answer.

OpenStudy (epicteatime):

oh its that the other vortex is also 1 and -1

OpenStudy (anonymous):

Let me check, calm down please.

OpenStudy (epicteatime):

thats why

OpenStudy (anonymous):

y = x^2 - 2(1) y = 1^2 - 2(1) y = 1 - 2 y = -1 Vertex for y = x^2 - 2x (1,-1)

OpenStudy (epicteatime):

oh so you got that too?

OpenStudy (anonymous):

Okay, so you are correct for the first equation. I'll check the second if you don't get upset.

OpenStudy (epicteatime):

*heavy breathing like a psycho

OpenStudy (anonymous):

*hands Epicteatime a chill pill*

OpenStudy (igreen):

@jcpd910 Here's the easy way out: https://www.desmos.com/calculator

OpenStudy (epicteatime):

*receives

OpenStudy (igreen):

Err @Epicteatime

OpenStudy (epicteatime):

NO

OpenStudy (anonymous):

iGreen, what about when he's doing a written exam and he doesn't have the option to use a quad calc?

OpenStudy (epicteatime):

lol

OpenStudy (anonymous):

He should know how to use this, so calculators are not the option. If they were, nobody would have to do real math.

OpenStudy (anonymous):

Epic, want me to check your work?

OpenStudy (igreen):

Actually, they would. Graphing calculators only make it easier to graph things.

OpenStudy (anonymous):

for the second eqaution?

Parth (parthkohli):

?

OpenStudy (epicteatime):

yep

OpenStudy (anonymous):

iGreen, it's nice to know how to do this. -.-

OpenStudy (epicteatime):

yes please

OpenStudy (anonymous):

Right, going in! *puts on war paint*

OpenStudy (igreen):

Lol..I already know the long way of doing it..so I just use the Calculator..it's just faster.

OpenStudy (epicteatime):

lol

OpenStudy (anonymous):

y = -2x^2 + 4x - 3 \[x = \frac{ -b \pm \sqrt{b ^{2} - 4ac }}{ 2a }\] a = -2 b = 4 c = -3 \[x = \frac{ -4 \pm \sqrt{4^{2} - 4(-2)(-3) }}{ 2(-2) }\] \[x = \frac{ -b \pm \sqrt{16 - 24 }}{ -4 }\] \[x = \frac{ -4 \pm \sqrt{-8}}{ -4 }\] hmmm looks like we're gonna have to use complex solutions, which I think means we aren't gonna have (1,-1)

OpenStudy (epicteatime):

WOOOOO

OpenStudy (epicteatime):

what the crap is this

OpenStudy (anonymous):

Wait though, a complex solution still is a solution, to the point where we can solve it. problem with me is that I'm not good at complex solutions. O.o

OpenStudy (epicteatime):

lol

OpenStudy (epicteatime):

so...

OpenStudy (anonymous):

Have you been taught complex solutions yet?

OpenStudy (epicteatime):

nope

OpenStudy (epicteatime):

i 'started' learning this like a few days ago...